A sample of 31 students in normally distributed with a mean age of 22.6 years and a standard deviation of 1.6 years. Construct a 95% confidence interval for the population standard deviation of student ages. Round the boundaries to two decimal places.
Answer #1

Soln n=31,7= 22.6 s=1.6 ,c=95 1,2=0.08 dr = n-1=30 Pa, 20, 0.025 = 46.979 the hot, i-am = 420,0975 = 16.791 scholl coast chil